OpenStudy (anonymous):

jolene paid $10.50 to bowl 3 games.She also paid $2.50 to rent bowling shoes. How much did she pay per game?

OpenStudy (anonymous):

3+3=6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3/10.50

OpenStudy (anonymous):

What does that mean

OpenStudy (anonymous):

3.5$

OpenStudy (anonymous):

the answer

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thankyou

OpenStudy (texaschic101):

(10.5/3) + 2.50 = 3.50 + 2.50 = $6

OpenStudy (anonymous):

Is it $6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how

OpenStudy (anonymous):